# A Systematic Survey and Critical Review on Evaluating Large Language Models: Challenges, Limitations, and Recommendations

### v1

@@ 1 Introduction
To mitigate this gap, this paper brings together the discussions to address the fundamental challenges and limitations in LLM evaluations that emerge from diverse evaluation setups. First, we craft a schematic workflow of the evaluation pipeline in practical settings (presented in Section 2) for a systematic study. We then examine each step in the evaluation workflow, uncovering various inconsistencies and decision-making complexities affecting reproducibility, reliability, and robustness (see Section 3). Based on our findings, we provide a principled guideline in Section 4 to address current limitations in LLM evaluation.
@@ 3.1.2 Lacking Response Generation Details
Prompting: The lack of details behind how the prompts are designed may make the findings in different literature inconsistent. For instace, variations in prompt design can lead to significantly different results, as seen in various studies Jahan et al. (2024); Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023). While few-shot learning is found to outperform zero-shot in the original evaluation conducted by the authors of various LLMs OpenAI (2023); Anil et al. (2023); Touvron et al. (2023b), many independent evaluations demonstrate that adding few-shot examples does not necessarily outperform zero-shot models in every task Ye et al. (2023a); Jahan et al. (2024). This raises the concern of whether certain prompt engineering techniques or optimizations to select few-shot samples were applied in the original evaluations. Hence, not disclosing the details behind how the prompt is designed or how the few-shot examples are selected can hinder reproducibility.
@@ 3.2.1 Data and Model Integrity Issues
Contamination in Existing Models: Contamination occurs when a benchmarking dataset is used in training, reducing result reliability and validity Sainz et al. (2023a); Zhou et al. (2023b); Shi et al. (2023). Ensuring benchmarking examples are excluded from training data is essential to maintain reliable results. Since LLMs are pre-trained on vast amounts of text data available on the internet, this could lead to unfair evaluations if LLMs have already encountered these datasets during their pre-training phase Ravaut et al. (2024); Xu et al. (2024); Balloccu et al. (2024).
Nonetheless, most prior LLM evaluation work focusing on zero-shot evaluation did not conduct any data contamination tests Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); OpenAI (2023); Team et al. (2023), raising concerns about whether these evaluations truly represent the zero-shot capabilities of LLMs. Recent research has also demonstrated a strong possibility of data contamination in many datasets used to evaluate different LLMs Sainz et al. (2023b); Li and Flanigan (2023); Golchin and Surdeanu (2023); Ravaut et al. (2024); Xu et al. (2024); Zhang et al. (2024a); Oren et al. (2023); Balloccu et al. (2024). With the current generation of LLMs being extremely capable of learning new skills with minimal amounts of data, exposing them to evaluation data may undermine the measurement of their true capabilities. Since the possibility of data contamination has led to the development of new versions of existing datasets (e.g., utilizing GSM-8K to construct GSM-1K Zhang et al. (2024a)), it is crucial to use reliable and fair evaluation datasets.
@@ 3.3 Robustness
While there are many evaluation benchmarks currently available, existing work mostly relies on evaluating LLMs on some common benchmarks, this raises the question of whether the performance of LLMs in these common benchmarks in existing settings reflects their true capabilities and limitations. In this section, we study the robustness of existing LLM evaluations.
@@ 3.3.1 Lacking Generalized Evaluation
Limiting Evaluation to Certain Scenarios: Interestingly, it has been observed in recent research that certain performance gains in a specific dataset may not necessarily imply that it would also improve the performance in other datasets for similar tasks Jahan et al. (2024); SambaNova (2024). For instance, Jahan et al. (2024) observes that not a single LLM has superiority over other LLMs across all biomedical datasets and tasks. This is also evident if we compare the results between LLaMA-3 and Qwen2 reported in Qwen2 (2024). As shown in Figure 3, while the Qwen2 model outperforms LLaMA-3 on most datasets, it falls short on GPQA and MBPP. Interestingly, for coding tasks, Qwen2 significantly outperforms LLaMA-3 on the HumanEval dataset Chen et al. (2021) but not on the MBPP dataset Austin et al. (2021). Meanwhile, existing common benchmarks also do not take into account some specific settings, such as how LLMs perform in long context scenarios, as recent research demonstrated that LLMs often struggle to generate the correct answer when relevant information does not appear at the beginning or end of the input context Liu et al. (2024c). This highlights the importance of evaluating the generalized performance of LLMs across a set of diverse benchmarks and settings,instead of limiting evaluation to only common benchmarks like MMLU Hendrycks et al. (2020b).
@@ 3.3.2 No Tuning of Prompt and Decoding Parameters
As shown in Figure 4, the restricted prompting method by Laskar et al. (2023a) consistently outperforms the unrestricted approach across all three models. Conversely, the restricted prompting method by Qin et al. (2023) fails to surpass the unrestricted approach for GPT-3.5 and GPT-4o. However, it surprisingly outperforms the unrestricted method, indicating the significant impact of prompt tuning across models. Evaluating language models with a single prompt lacks fairness Zhu et al. (2023b), yet it remains common practice Bang et al. (2023); Qin et al. (2023); Laskar et al. (2023a). Minor prompt variations can lead to diverse outcomes for different models Alzahrani et al. (2024); An et al. (2023); Zhang et al. (2024a); Sclar et al. (2023); Lanham et al. (2023); Biderman et al. (2024), highlighting the need to compare benchmarks across multiple prompts. Using automated prompt tuning techniques like Meta Probing Agents Zhu et al. (2024) can ensure robustness to prompt variations.
@@ 4 Recommendations and Best Practices
We’ve outlined the primary challenges in evaluating LLMs thus far. In light of these challenges, a crucial question arises: How can we enhance the evaluation process for LLMs? Crafting a structured framework that’s both practical and easy to implement is daunting, given the complexities of generative LLM development. Previous studies have tended to focus on specific evaluation aspects without offering comprehensive guidelines for the entire evaluation cycle, leaving researchers without clear guidance. Before diving into recommendations for each evaluation stage, it’s important to acknowledge three key factors shaping current LLM evaluation practices: inherent randomness in generative models, significant computational demands, and insufficient documentation across stages.
@@
Evaluation Setup: Selecting benchmarks for model assessment is crucial. Rather than simply replicating past choices, researchers should align datasets with required capabilities. To ensure robustness, datasets should vary across expected LLM capabilities (e.g., long-context understanding), tasks (e.g., summarization), and language complexity (e.g., vocabulary coverage). Ideally, a metric should measure dataset diversity. For model selection, conduct contamination tests between the chosen model and benchmarks using relevant techniques Ravaut et al. (2024). This acts as an additional filter for benchmarking datasets, ensuring selection of unseen ones measuring intended capabilities. Meanwhile, for reproducibility, document any subset use of benchmarking datasets, along with the selected model version. In addition, throughout scientific history, intelligence progress has evolved across generations. Tests from a decade ago may appear simplistic compared to today’s standards (e.g., Math Olympiads, ICPC programming contests). Refreshing LLM evaluations periodically can effectively communicate standard capabilities in both open and closed-source LLM markets and ecosystems (e.g., chatbots, translation tools). Hence, to ensure reliability, verify if the dataset has updated versions and incorporate them if available (e.g., HumanEvalPlus Liu et al. (2024b), MMLU-Pro Wang et al. (2024b))
@@
Response Generation: For reproducibility, thorough documentation of prompts and parameter settings is essential (e.g., specifying how few-shot samples are selected). To ensure reliability, it’s crucial to justify why specific prompts and parameter settings are chosen over others, and provide comparisons with alternative options. As for robustness, running experiments with diverse prompts and parameters is key to showcasing their effectiveness and limitations across different scenarios. In resource-constrained environments, conducting experiments with diverse evaluation settings may pose challenges, yet it remains vital to perform robust evaluations on at least a subset of samples.
@@
Evaluation Methodology: To ensure reproducibility, the parsing scripts and the output data used for evaluation should be published. Meanwhile, sanity-checking on the parsing script should be done to ensure reliability and robustness of the designed parsing script. This can be done by creating test cases for various response types, and then verifying (with human intervention if possible) whether the parsing script can reliably extract the targeted answer from the generated response. Meanwhile, reliance on string-based metrics like ROUGE should be minimized in favor of qualitative evaluations to ensure the reliability of the chosen evaluation methodology. Given the cost and time constraints of human qualitative evaluation, LLM-based evaluators can be used as alternatives but must be validated for potential biases (e.g., multiple LLMs as juries instead of using a single LLM as the judge Zheng et al. (2024)). Finally, robust evaluation using task-specific metrics is encouraged with the metrics that lack alignments with humans should be avoided.
@@ 5 Conclusions and Future Work
In this paper, we systematically survey the challenges and limitations in evaluating LLMs. We identified significant inconsistencies and complexities at various stages of the evaluation pipeline, impacting the reproducibility, reliability, and robustness of the results. These issues underline the necessity for a standardized and systematic approach to evaluating LLMs to ensure their reliable usage in real-world applications. By comprehensively reviewing the current evaluation practices, we have provided a set of recommendations aimed at enhancing the consistency and reliability of LLM evaluations. Therefore, future work should focus on developing and adopting standardized evaluation protocols for LLMs to address the identified inconsistencies and complexities. This includes creating benchmark datasets, evaluation metrics, and proper documentation of the evaluation settings to ensure reproducibility, reliability, and robustness.
@@ A.1 Benchmarking Datasets
There are also specialized benchmarks that measure specific capabilities of LLMs. For instance, the MT-Bench Zheng et al. (2024)) evaluates whether LLMs can properly engage in conversations, the RewardBench Lambert et al. (2024) assesses the performance of reward models. Other specialized benchmarks like the AlpacaEval777https://tatsu-lab.github.io/alpaca_eval/ evaluates the instruction following capabilities Zhou et al. (2023a) of LLMs, the Open Medical-LLM Leaderboard888https://huggingface.co/blog/leaderboard-medicalllm evaluates the biomedical capabilities of LLMs, HHEM999https://huggingface.co/spaces/vectara/leaderboard leaderboard for hallucination detection Mishra et al. (2024); Sadat et al. (2023), BigCodeBench101010https://huggingface.co/blog/leaderboard-bigcodebench and LiveCodeBench111111https://huggingface.co/blog/leaderboard-livecodebench for code generation capability evaluation, SWE-bench Jimenez et al. (2023) for software engineering capability evaluation. The recently proposed FOFO benchmark Xia et al. (2024) measures language models’ ability to adhere to the requested formats in prompts across different domains. Moreover, there are also some specialized benchmarks that are used for LLM safety121212https://huggingface.co/spaces/AI-Secure/llm-trustworthy-leaderboard Chao et al. (2024) and red teaming131313https://huggingface.co/spaces/HaizeLabs/red-teaming-resistance-benchmark Tedeschi et al. (2024) evaluation. The ability to understand both language and vision is another interesting capability of recently proposed LLMs Zhu et al. (2023a); Liu et al. (2024a, 2023b); Dai et al. (2024); Zhang et al. (2023); Ye et al. (2023b); Luo et al. (2024); Bai et al. (2023); Chen et al. (2023a). This has led to the development of many multi-modal benchmarks Fu et al. (2023a); Yu et al. (2023); Liu et al. (2023e); Li et al. (2023a, b); Liu et al. (2024a); Fu et al. (2024a); Lu et al. (2022); Guan et al. (2023); Li et al. (2023d); Qiu et al. (2024); Chen et al. (2024b). These benchmarks study the multimodal capabilities of LLMs across various domains, such as math and reasoning Yue et al. (2023); Lu et al. (2023), science diagrams Kembhavi et al. (2016), chart understanding and reasoning Masry et al. (2022); Islam et al. (2024b); Masry et al. (2024); Rahman et al. (2023), document understanding Mathew et al. (2021).
@@
To enable a more comprehensive evaluation of LLMs across a wide range of scenarios, some studies also focused on introducing new benchmarks covering various aspects, such as HELM Liang et al. (2022), PromptBench Zhu et al. (2023b), OpenLLM141414https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, MixEval Ni et al. (2024), etc. These benchmarks cover diverse tasks and usually include existing benchmarking datasets (e.g., MMLU, HellaSwag, BoolQ Clark et al. (2019), etc.). Additionally, despite the availability of numerous benchmarks (both general and specialized), existing widely-used benchmarks still do not cover the full variety of tasks Preum et al. (2018); Parvez et al. (2018). Therefore, some researchers have independently evaluated LLMs using additional diverse datasets and tasks, including various NLP datasets and tasks Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); Kocoń et al. (2023). They also employed domain-specific benchmarks in fields such as biomedicine Jahan et al. (2023, 2024), finance Li et al. (2023e); Guo et al. (2023a), language-specific Kabir et al. (2023); Lai et al. (2023); Khondaker et al. (2023); Ahuja et al. (2023); Liu et al. (2023c); Abdelali et al. (2024), social science Ziems et al. (2024), coding Liu et al. (2024b), and information retrieval Zhu et al. (2023c). In addition to that, ethics, bias, toxicity, robustness, and trustworthiness are also independently evaluated by researchers across various datasets Hendrycks et al. (2020a); Yang et al. (2022); Wang et al. (2023a); Zhuo et al. (2023); Rawte et al. (2023); Liu et al. (2023a); McIntosh et al. (2024); Sun et al. (2024).

### v2

@@ 1 Introduction
To mitigate this gap, this paper brings together the discussions to address the fundamental challenges and limitations in LLM evaluations that emerge from diverse evaluation setups. First, we craft a schematic workflow of the evaluation pipeline in practical settings (presented in Section 2) for a systematic study. We then examine each step in the evaluation workflow, uncovering various inconsistencies and decision-making complexities affecting reproducibility, reliability, and robustness (see Section 3). Based on our findings, we provide a principled guideline in Section 4 to address current limitations in LLM evaluation. The data and the code used in this paper are publicly available here: https://github.com/ntunlp/Critical-Review-of-LLM-Eval.
@@ 3.1.2 Lacking Response Generation Details
Prompting: The lack of details behind how the prompts are designed may make the findings in different literature inconsistent. For instance, variations in prompt design can lead to significantly different results, as seen in various studies Jahan et al. (2024); Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023). While few-shot learning is found to outperform zero-shot in the original evaluation conducted by the authors of various LLMs OpenAI (2023); Anil et al. (2023); Touvron et al. (2023b), many independent evaluations demonstrate that adding few-shot examples does not necessarily outperform zero-shot models in every task Ye et al. (2023a); Jahan et al. (2024). This raises the concern of whether certain prompt engineering techniques or optimizations to select few-shot samples were applied in the original evaluations. Hence, not disclosing the details behind how the prompt is designed or how the few-shot examples are selected can hinder reproducibility.
@@ 3.2.1 Data and Model Integrity Issues
Contamination in Existing Models: Contamination occurs when a benchmarking dataset is used in training, reducing result reliability and validity Sainz et al. (2023); Zhou et al. (2023b); Shi et al. (2023). Ensuring benchmarking examples are excluded from training data is essential to maintain reliable results. Since LLMs are pre-trained on vast amounts of text data available on the internet, this could lead to unfair evaluations if LLMs have already encountered these datasets during their pre-training phase Ravaut et al. (2024); Xu et al. (2024); Balloccu et al. (2024).
Nonetheless, most prior LLM evaluation work focusing on zero-shot evaluation did not conduct any data contamination tests Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); OpenAI (2023); Team et al. (2023), raising concerns about whether these evaluations truly represent the zero-shot capabilities of LLMs. Recent research has also demonstrated a strong possibility of data contamination in many datasets used to evaluate different LLMs Sainz et al. (2023); Li and Flanigan (2023); Golchin and Surdeanu (2023); Ravaut et al. (2024); Xu et al. (2024); Zhang et al. (2024a); Oren et al. (2023); Balloccu et al. (2024); Matton et al. (2024). With the current generation of LLMs being extremely capable of learning new skills with minimal amounts of data, exposing them to evaluation data may undermine the measurement of their true capabilities. Since the possibility of data contamination has led to the development of new versions of existing datasets (e.g., utilizing GSM-8K to construct GSM-1K Zhang et al. (2024a)), it is crucial to use fair evaluation datasets.
@@ 3.3 Robustness
In the context of evaluating LLMs, robustness refers to the model’s ability to maintain consistent performance across a wide range of inputs, conditions, or tasks. While there are many evaluation benchmarks currently available, existing work mostly relies on evaluating LLMs on some common benchmarks. This raises the question of whether the performance of LLMs in these common benchmarks reflects their true capabilities and limitations. In this section, we study the robustness of existing LLM evaluations.
@@ 3.3.1 Lacking Generalized Evaluation
Limiting Evaluation to Certain Scenarios: Interestingly, it has been observed in recent research that certain performance gains in a specific dataset may not necessarily imply that it would also improve the performance in other datasets for similar tasks Jahan et al. (2024); SambaNova (2024). For instance, Jahan et al. (2024) observes that not a single LLM has superiority over other LLMs across all biomedical datasets and tasks. This is also evident if we compare the results between LLaMA-3 and Qwen2 reported in Yang et al. (2024); Qwen2 (2024). As shown in Figure 3, while the Qwen2 model outperforms LLaMA-3 on most datasets, it falls short on GPQA and MBPP. Interestingly, for coding tasks, Qwen2 significantly outperforms LLaMA-3 on the HumanEval dataset Chen et al. (2021) but not on the MBPP dataset Austin et al. (2021). Meanwhile, existing common benchmarks also do not take into account some specific settings, such as how LLMs perform in long context scenarios, as recent research demonstrated that LLMs often struggle to generate the correct answer when relevant information does not appear at the beginning or end of the input context Liu et al. (2024c). This highlights the importance of evaluating the generalized performance of LLMs across a set of diverse benchmarks and settings,instead of limiting evaluation to only common benchmarks like MMLU Hendrycks et al. (2020b).
@@ 3.3.2 No Tuning of Prompt and Decoding Parameters
As shown in Figure 4, the restricted prompting method by Laskar et al. (2023a) consistently outperforms the unrestricted approach across all three models. Conversely, the restricted prompting method by Qin et al. (2023) fails to surpass the unrestricted approach for GPT-3.5 and GPT-4o. However, it surprisingly outperforms the unrestricted method, indicating the significant impact of prompt tuning across models. Evaluating language models with a single prompt lacks fairness Zhu et al. (2023b), yet it remains common practice Bang et al. (2023); Qin et al. (2023); Laskar et al. (2023a). Minor prompt variations can lead to diverse outcomes for different models Alzahrani et al. (2024); An et al. (2023); Zhang et al. (2024a); Sclar et al. (2023); Lanham et al. (2023); Biderman et al. (2024); Wei et al. (2024), highlighting the need to compare benchmarks across multiple prompts. Using automated prompt tuning techniques like Meta Probing Agents Zhu et al. (2024) can ensure robustness to prompt variations.
@@ 3.3.3 Evaluation Method’s Generalizability and Correlation Shortcomings
Step | Sub-Step | Recommendation | Implementation: Suggested Tools or Techniques |
Evaluation Setup | Benchmark Selection | Selected benchmarks should align with the capabilities required and updated versions of the datasets should be used to ensure reliability, diversity in the selected benchmarks is required to ensure robustness, and proper documentation of the dataset subsets is required for reproducibility. |
Reliability: Use refined benchmarks like MMLU-Pro, Human-Eval Plus, GSM-1k to address the limitations in existing benchmarks to improve reliability.
Reproducibility: Document the data sampling technique and release the data subset used for evaluation alongside the model-generated response. Robustness: Check tokenizer vocabulary coverage in selected benchmarks. |
Model Selection | Data contamination check in the selected model is required for reliability, proper versioning of the model is required for reproducibility, and diverse capability evaluation (e.g., latency, memory usage, format following capability, etc.) is important to ensure robustness. |
Reliability: Use tools like LLMSanitize Library Ravaut et al. (2024) for contamination check.
Reproducibility: Use mlflow or W&B for documentation. Robustness: Use tools like pyNVML to measure GPU memory requirements, FOFO for format following, compare accuracy vs latency trade-off, etc. |
Response Generation | Prompt Design | Release the prompts and few-shot examples for reproducibility, justify the selection of certain prompts and few-shot examples to ensure reliability, and compare with alternative prompts to ensure robustness. |
Reliability: Justify the choice of certain prompts to ensure no potential of prompt hacking and compare the alternatives. Meanwhile, clearly demonstrate what and how few-shot examples are selected.
Reproducibility: Use tools like LM-Evaluation-Harness. Robustness: Use Prompt Bench or Meta-Probing Agent. |
Decoding Parameters | Document the decoding parameters to ensure reproducibility, justify the selection to ensure reliability, and experiment with various parameters to ensure robustness. |
Reliability: Justify the choice of certain parameters to eliminate the risk of optimization in the test data.
Reproducibility: Use mlflow or W&B. Robustness: Compare the performance based on different decoding parameters, at least in a subset of the data. |
Evaluation Methodology | Parsing Script Design | Accurate parsing of the response is required for reliability, availability of these scripts is needed for reproducibility, and parsing scripts should show robustness across different models and datasets. |
Reliability: Validate the reliability based on human evaluation, at least on a subset.
Reproducibility: Release the code. Robustness: Evaluate multiple models and datasets, across all types of labels and corner cases. |
Evaluation Approach | Availability of the evaluation output is required for reproducibility, selected evaluation metrics should maintain correlation with humans to ensure reliability, and multiple evaluation metrics are required for evaluation robustness. |
Reliability: Validate the effectiveness of selected metrics (e.g., measure correlation with humans), use techniques like LLM-as-juries to mitigate bias.
Reproducibility: Release the Evaluation Output. Robustness: Use multiple evaluation metrics (e.g., in Summarization, use both word-based (e.g., ROUGE) or Contextualized (e.g., BERTScore) metrics), measure latency, GPU usage via pyNVML. |
@@ 4 Recommendations and Best Practices
So far, we’ve outlined the primary challenges in evaluating LLMs. In light of these challenges, a crucial question arises: How can we enhance the evaluation of LLMs? Crafting a structured framework that’s both practical and easy to implement is daunting, given the complexities of generative LLM development. Previous studies tended to focus on specific evaluation aspects without offering comprehensive guidelines for the entire evaluation cycle, leaving researchers without clear guidance. Before diving into recommendations for each evaluation stage, it’s important to acknowledge three key factors shaping current LLM evaluation practices: inherent randomness in generative models, significant computational demands, and insufficient documentation across stages.
@@
Evaluation Setup: Selecting benchmarks for model assessment is crucial. Rather than simply replicating past choices, researchers should align datasets with required capabilities. To ensure robustness, datasets should vary across expected LLM capabilities (e.g., long-context understanding), tasks (e.g., summarization), and language complexity (e.g., vocabulary coverage). Ideally, a metric should measure dataset diversity. For model selection, conduct contamination tests between the chosen model and benchmarks using relevant techniques Ravaut et al. (2024). This acts as an additional filter for benchmarking datasets, ensuring selection of unseen ones measuring intended capabilities. Meanwhile, for reproducibility, document any subset use of benchmarking datasets, along with the selected model version. In addition, throughout scientific history, intelligence progress has evolved across generations. Tests from a decade ago may appear simplistic compared to today’s standards (e.g., Math Olympiads, ICPC programming contests). Refreshing LLM evaluations periodically can effectively communicate standard capabilities in both open and closed-source LLM markets and ecosystems (e.g., chatbots). Hence, to ensure reliability, verify if the dataset has updated versions and incorporate them if available (e.g., HumanEvalPlus Liu et al. (2024b), MMLU-Pro Wang et al. (2024b), GSM-1K Zhang et al. (2024a))
@@
Response Generation: For reproducibility, thorough documentation of prompts (e.g., explaining the selection of few-shot samples) and parameter settings (e.g., use tools like mlflow777https://mlflow.org/ or Weights & Biases888https://wandb.ai/site (W&B)) is essential. To ensure reliability, it’s crucial to justify why specific prompts and parameters are chosen over others by providing comparisons with alternative options. As for robustness, experimenting with diverse prompts and parameters is the key to showcasing their effectiveness and limitations in different scenarios. In resource-constrained environments, conducting experiments with diverse evaluation settings may pose challenges, yet it remains vital to perform robust evaluations on at least a subset of samples.
@@
Evaluation Methodology: To ensure reproducibility, the parsing scripts and the output data used for evaluation should be published. Meanwhile, sanity-checking on the parsing script should be done to ensure reliability and robustness of the designed parsing script. This can be done by creating test cases for various response types, and then verifying (with human intervention if possible) whether the parsing script can reliably extract the targeted answer from the generated response. Meanwhile, reliance on string-based metrics like ROUGE should be minimized in favor of qualitative evaluations to ensure the reliability of the chosen evaluation methodology. Given the cost and time constraints of human qualitative evaluation, LLM-based evaluators can be used as alternatives but must be validated for potential biases (e.g., multiple LLMs as juries instead of using a single LLM as the judge Zheng et al. (2024)). Finally, robust evaluation using task-specific metrics is encouraged. For this purpose, metrics that lack alignment with humans should be avoided. Moreover, measuring runtime latency using tools like pyNVML999https://pypi.org/project/pynvml/ is recommended to evaluate the real-world applicability of different LLMs.
@@ 5 Conclusions and Future Work
In this paper, we systematically survey the challenges and limitations in evaluating LLMs. We identified significant inconsistencies and complexities at various stages of the evaluation pipeline, impacting the reproducibility, reliability, and robustness of the results. These issues underline the necessity for a standardized and systematic approach for LLM evaluation to ensure their reliable usage in real-world applications. By comprehensively reviewing the current evaluation practices, we have provided a set of recommendations aimed at enhancing the consistency and fairness of LLM evaluations. Therefore, future work should focus on developing and adopting standardized evaluation protocols for LLMs to address the identified limitations. This includes creating benchmark datasets, evaluation metrics, and proper documentation of the evaluation settings to ensure reproducibility, reliability, and robustness.
Acknowledgements
We would like to thank all the anonymous reviewers for their excellent review comments. This research was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the York Research Chairs (YRC) program. We also acknowledge Compute Canada for the computing resources. Finally, we thank Mir Tafseer Nayeem for providing valuable feedback.
@@ Acknowledgements
Moreover, given the limited amount of study to evaluate LLMs in non-English data, our work was more focused on the monolingual scenario (mostly on English data). Therefore, investigating the challenges and limitations of LLM evaluation in multilingual and resource-constrained scenarios could be studied in the future, alongside also studying the performance of various tokenizers (both multilingual and monolingual) in LLM benchmarking Rust et al. (2021); Choo and Kim (2023)).
@@ Appendix A Appendix
Criteria | Challenges & Limitations | Description |
Reproducibility | Missing Experimental Details | Lack of documentation on the data subsets used for evaluation, which few-shot examples added to the prompt, what decoding parameters are used, etc., will impact reproducibility. |
Not Releasing the Data | The detailed prompt as well as the response generated by the LLMs are often missing. | |
Code Unavailable | Many studies do not release the necessary codes (e.g., parsing scripts). This may impact reproducibility of the results. | |
Model Updates and Depreciation | Continuous updates to the closed-source models (alongside possible depreciation of the models) will create challenges for reproducing previous results. | |
Reliability | Not Documenting Model Versions | The exact version of the model being used is often missing. This creates another reproducibility concern. |
Data Integrity | Incorrect gold labels and outdated benchmark datasets compromise evaluation reliability. | |
Unfair Comparisons | Comparing models evaluated on the full dataset against the subset of a dataset, different few-shot examples being selected, etc. | |
Contamination | LLMs may encounter evaluation data during pre-training, leading to contamination. | |
Prompt Hacking | Manipulating input prompts to elicit desired responses can undermine fair evaluation. | |
Transparency in Decoding Parameters | Lack of transparency in how decoding parameters are selected can lead to unfair comparisons. | |
Robustness | Evaluation Methodology and Metrics | Reliance on string-based metrics and automated evaluation methods without proper validation can lead to unreliable results. |
Limiting Evaluation to Certain
Benchmarks |
Evaluating LLMs only on a set of common benchmarks does not ensure generalizability. | |
Lack of Diversity in Prompts
and Parameters |
Most existing research used only a single prompt while also not tuning any of the decoding parameters, restricting the robustness of the evaluation. | |
Insufficient Evaluation Metrics | Lack of correlation between existing evaluation metrics impacts evaluation robustness. |
@@ A.1 Benchmarking Datasets
There are also specialized benchmarks that measure specific capabilities of LLMs. For instance, the MT-Bench Zheng et al. (2024)) evaluates whether LLMs can properly engage in conversations, the RewardBench Lambert et al. (2024) assesses the performance of reward models. Other specialized benchmarks like the AlpacaEval101010https://tatsu-lab.github.io/alpaca_eval/ evaluates the instruction following capabilities Zhou et al. (2023a) of LLMs, the Open Medical-LLM Leaderboard111111https://huggingface.co/blog/leaderboard-medicalllm evaluates the biomedical capabilities of LLMs, HHEM121212https://huggingface.co/spaces/vectara/leaderboard leaderboard for hallucination detection Mishra et al. (2024); Sadat et al. (2023), BigCodeBench131313https://huggingface.co/blog/leaderboard-bigcodebench and LiveCodeBench141414https://huggingface.co/blog/leaderboard-livecodebench for code generation capability evaluation, SWE-bench Jimenez et al. (2023) for software engineering capability evaluation. The recently proposed FOFO benchmark Xia et al. (2024) measures language models’ ability to adhere to the requested formats in prompts across different domains. Moreover, there are also some specialized benchmarks that are used for LLM safety151515https://huggingface.co/spaces/AI-Secure/llm-trustworthy-leaderboard Chao et al. (2024) and red teaming161616https://huggingface.co/spaces/HaizeLabs/red-teaming-resistance-benchmark Tedeschi et al. (2024) evaluation. The ability to understand both language and vision is another interesting capability of recently proposed LLMs Zhu et al. (2023a); Liu et al. (2024a, 2023b); Dai et al. (2024); Zhang et al. (2023); Ye et al. (2023b); Luo et al. (2024); Bai et al. (2023); Chen et al. (2023a). This has led to the development of many multi-modal benchmarks Fu et al. (2023a); Yu et al. (2023); Liu et al. (2023e); Li et al. (2023a, b); Liu et al. (2024a); Fu et al. (2024a); Lu et al. (2022); Guan et al. (2023); Li et al. (2023d); Qiu et al. (2024); Chen et al. (2024b). These benchmarks study the multimodal capabilities of LLMs across various domains, such as math and reasoning Yue et al. (2023); Lu et al. (2023), science diagrams Kembhavi et al. (2016), chart understanding and reasoning Masry et al. (2022); Islam et al. (2024b); Masry et al. (2024); Rahman et al. (2023), document understanding Mathew et al. (2021).
@@
To enable a more comprehensive evaluation of LLMs across a wide range of scenarios, some studies also focused on introducing new benchmarks covering various aspects, such as HELM Liang et al. (2022), PromptBench Zhu et al. (2023b), OpenLLM171717https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, MixEval Ni et al. (2024), etc. These benchmarks cover diverse tasks and usually include existing benchmarking datasets (e.g., MMLU, HellaSwag, BoolQ Clark et al. (2019), etc.). Additionally, despite the availability of numerous benchmarks (both general and specialized), existing widely-used benchmarks still do not cover the full variety of tasks Preum et al. (2018); Parvez et al. (2018). Therefore, some researchers have independently evaluated LLMs using additional diverse NLP datasets and tasks Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); Kocoń et al. (2023). They also employed domain-specific benchmarks in fields such as biomedicine Jahan et al. (2023, 2024), finance Li et al. (2023e); Guo et al. (2023a), language-specific Kabir et al. (2023); Lai et al. (2023); Khondaker et al. (2023); Ahuja et al. (2023); Liu et al. (2023c); Abdelali et al. (2024), social science Ziems et al. (2024), coding Liu et al. (2024b), and information retrieval Zhu et al. (2023c). In addition to that, ethics, bias, toxicity, robustness, and trustworthiness are also independently evaluated by researchers across various datasets Hendrycks et al. (2020a); Yang et al. (2022); Wang et al. (2023a); Zhuo et al. (2023); Rawte et al. (2023); Liu et al. (2023a); McIntosh et al. (2024); Sun et al. (2024).